OpenStudy (anonymous):

Write a recursive formula for the sequence 15, 26, 48, 92, 180, ... Then find the next term.

OpenStudy (anonymous):

15*2-4

OpenStudy (anonymous):

26*2-4

OpenStudy (anonymous):

180*2-4

OpenStudy (anonymous):

What's a recursive formula tho?

OpenStudy (anonymous):

Sargams is right, but you can also write: 26=15+(2^0)*11 48=26+(2^1)*11 92=48+(2^2)*11 180=92+(2^3)*11 So, in general, f(n+1)=f(n)+( 2^(n) )*11, starting with n=0 and f(0)=15.

OpenStudy (anonymous):

a formula which repeats itself... term*2-4=successive term

OpenStudy (anonymous):

I don't have an answer choice like that tho \(a_n= 2a_n – 1 – 4,~ where~ a1 = 15; 356\) \(a_n= 2a_n – 4,~ where~ a1 = 15; 356\) \(a_n= 4 + 11 · 2^{n–1},~ where~ a1 = 15; 356\) \(a_n= 3a_n – 1 – 19,~ where~ a1 = 15; 356\)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Write that one for your teacher and he/she is going to like it. xD